CHUCK INVESTED $14000 LAST YEAR. HE INVESTED 9% IN A SIMPLE INTEREST, AND THE REST OF IT AT A 10%. THEN MADE A TOTAL OF $1350 FROM INTEREST. HOW MUCH WOULD HAVE BEEN INESTED AT EACH RATE? WOULD ENJOY ANY HELP GIVEN. KEVIN
Answer by Nate(3500)

x(0.09) + (14000 - x)(.1) = 1350
-0.01x = -50
x = 5000
$5000 went into 9% while $9000 went into 10%